OpenStudy (anonymous):

Gold in the form of solid cylinders will be molded into solid blocks. Each cylinder has a diameter of 6.18 cm and a height of 6 cm. Each solid block will have a volume of 360 cm3. How many cylinders of gold will it take to mold one block of gold?

OpenStudy (anonymous):

the formula for working out volume of cylinders is \[\pi \times r ^{2}\times h\]

OpenStudy (anonymous):

you work out how much gold is in each cylinder first

OpenStudy (anonymous):

116.43?

OpenStudy (anonymous):

Would it be 3?